OpenStudy (anonymous):

The length of a football field is 67 yards greater than the width. The perimeter of the field is 346 yards. Find the length and width of the football field.

OpenStudy (anonymous):

Lenght times Width easy as that

OpenStudy (anonymous):

how do i find the width?

OpenStudy (anonymous):

width is 53 and length is 120

OpenStudy (anonymous):

so i would take the 346 and divide by 67 and that would give me the length?

OpenStudy (anonymous):

no i will show you to width

geerky42 (geerky42):

We are given that L = 67+W and P = 346. Knowing that P = 2L + 2W, we can replace L to (67 + W) and P to 346, like this: P = 2L + 2W 346 = 2(67+W) + 2W Can you find W?

OpenStudy (anonymous):

take the 53 and add 67 and that would make it 120?

OpenStudy (anonymous):

yes! i see it! :D

geerky42 (geerky42):

Where did you get 53?

OpenStudy (anonymous):

let width is x ok

OpenStudy (anonymous):

therfor length is x+67

OpenStudy (anonymous):

as we know perimeter = 2(l+b)

OpenStudy (anonymous):

put values of length and width here

OpenStudy (anonymous):

so 2(x+67) +2x

OpenStudy (anonymous):

=346

geerky42 (geerky42):

yes

OpenStudy (anonymous):

346= 2(x+x+67) 346=4x+134

geerky42 (geerky42):

then solve for x.

OpenStudy (anonymous):

so the length is 134 and the width is 53. because when i solved for x i got 53

OpenStudy (anonymous):

346-134=4x 212=4x 212/4=x 53=x

OpenStudy (anonymous):

i tried to check my answer with 2L+2W and i got 374 though?

OpenStudy (anonymous):

recheak

OpenStudy (anonymous):

134 x 2=268 53 x 2= 106 106+268=374

OpenStudy (anonymous):

53+67 = 120 not 134

OpenStudy (anonymous):

oh x) i had to go back and look at the work i see it now 120 x 2=240 53 x 2=106 240+106=346